OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today? 900 620 520 450

OpenStudy (anonymous):

Let x = one-step and y = two-step. x + y = 1070 3x - 2y = 110 Are you able to solve this simulataneous system?

OpenStudy (anonymous):

2x + 2y = 2140 3x - 110 = 2y 2x + (3x - 110) = 2140 5x - 110 = 2140 5x = 2250 x = 450 y = 620

OpenStudy (anonymous):

And that's all there is to it @Solmyr

OpenStudy (anonymous):

thanks again